# Partial identification of principal causal effects under violations of principal ignorability

### v2

@@ Abstract
Principal stratification,
Principal ignorability,
Partial identification,
@@ 1 Introduction
Another line of research aims to incorporate parametric assumptions to obtain identification when nonparametric identification is difficult or impossible to obtain. This is particularly the case with continuous intermediate variables, which we address in this manuscript, as standard assumptions such as monotonicity and exclusion restrictions are not sufficient for nonparametric identification. Recent works studying nonparametric identification in this more difficult scenario have made the principal ignorability assumption, along with the strong parametric assumption that the joint distribution of the two potential intermediates is governed by a known copula with a known correlation parameter [19, 28, 42]. Results are then estimated across a range of plausible correlation parameters to see if results are robust to this choice. Other works have attempted to estimate this unknown correlation parameter by jointly modeling the potential intermediates and observed outcomes simultaneously [37, 4]. The idea behind this strategy is that one can’t directly estimate the association between the two potential intermediates because they are never jointly observed, but this information can be obtained from a correctly specified outcome regression model. Related research aims to jointly model the distribution of the potential intermediates and observed outcome using either parametric [21] or semiparametric models [24, 25], and assume the correlation between the two potential intermediates is known. Many of these model-based approaches do not make the principal ignorability assumption, and they allow the observed outcome to depend on the values of both potential intermediate values.
@@ 2.2 Review of identification assumptions in PS analysis
$10<P(T=1\mid\boldsymbol{X}=\boldsymbol{x})<$ 1 for all $\boldsymbol{x}$.
@@ 2.2.1 Binary intermediates
While plausible in certain settings, such as those with randomized treatment assignment and noncompliance, these are frequently not likely to hold, as there are many settings where a direct effect of treatment is plausible. In such settings, one can instead use the following assumption
@@
2.2.2 Continuous intermediates
For a binary $S$, there are only four principal strata, and the monotonicity assumption eliminates one of these groups and enables the identification of the probability mass of $\boldsymbol{U}\mid\boldsymbol{X}$ [1]. For a continuous $S$, there are infinitely many principal strata, and while the monotonicity assumption can rule out some of them, there still remain infinitely many principal strata. Thus, monotonicity alone cannot produce the identification of the distribution of $\boldsymbol{U}\mid\boldsymbol{X}$, sometimes referred to as the principal density for continuous intermediates. Typically, researchers make parametric assumptions in order to identify this distribution. One approach is to assume the joint distribution of principal strata follows a known copula function $\mathbb{C}_{\rho}(\cdot,\cdot)$ (see [28]; [20]; [4]):
$pr(S(0)\leq s_{0},S(1)\leq s_{1}\mid\boldsymbol{X})=\mathbb{C}_{\rho}(pr(S(0)\leq s_{0}\mid\boldsymbol{X}),pr(S(1)\leq s_{1}\mid\boldsymbol{X})),$
@@ 3.1 A simple parametric model
$\begin{split}Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t&\sim N(\mu_{yit},\sigma_{y}^{2}),\qquad t=0,1.\\ \boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}&\sim N(\boldsymbol{\mu}_{si},\boldsymbol{\Sigma}_{s}),\end{split}$ (1)
@@
$\boldsymbol{\Sigma}_{s}=\begin{pmatrix}\sigma_{s0}^{2}&\rho\sigma_{s0}\sigma_{s1}\\ \rho\sigma_{s0}\sigma_{s1}&\sigma_{s1}^{2}\end{pmatrix}.$
@@ 3.2 Identification of the association parameter of the principal strata
where $\bar{T}=\sum\limits_{i=1}^{n}T_{i}/n$, $\zeta^{*}_{0}-{\psi^{*}_{0}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}$, and $\zeta_{1}^{*}-{\psi^{*}_{1}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}$.
@@ 4.1 Partial identification and principal causal effects
where $\mu_{yi}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{x},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{x})+\boldsymbol{\gamma}^{T}\boldsymbol{x}$, $\mu_{si}=\phi_{t}+\boldsymbol{\alpha}^{T}\boldsymbol{x}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$ for $t=0,1$. Clearly the parameters of the observed data distribution are identifiable from the observed data, and given $\boldsymbol{X}_{i}=\boldsymbol{x}$, there are 5 identifiable parameters for each of $t=0,1$ leading to 10 identifiable parameters. Excluding the parameters corresponding to $\boldsymbol{X}$ given by $(\boldsymbol{\gamma},\boldsymbol{\alpha})$, Model (1) has 11 parameters, showing that there are fewer known equations than unknown variables. This leads to partial identification of certain parameters, despite being the simplest, parametric model possible that allows for violations of principal ignorability. Naturally, this also implies that the parameters of more complex semi- or nonparametric models, such as those seen in the literature [4, 37], are at best partially identified in this setting unless additional constraints are imposed on the model parameters.
@@ 5 Weaker alternative assumptions of principal ignorability
For $\beta_{00}$ and $\beta_{11}$ with signs different from those assumed in Proposition 2, the corresponding partial identification regions for $\beta_{01}$ and $\beta_{10}$ can be obtained by flipping the signs to match those of $\beta_{00}$ and $\beta_{11}$, respectively. This result also requires us to know the sign of $\beta_{tt}$ for $t=0,1$. If this is not known a priori, then one can regress the observed $Y$ on the observed $S$ and $\boldsymbol{X}$, given $T=t$, to determine the sign for $\beta_{tt}$. This approach is not guaranteed to correctly identify the sign, but will only fail to do so in certain extreme situations, such as when $\rho$ is close to $-1$. This cuts the partial identification region in half, providing large benefits in the uncertainty in the resulting causal estimates. While these benefits may be large, it is not possible in many situations to know this assumption is true a priori. Now we present a second, weaker assumption, which should hold true in most applications. This second assumption effectively assumes that the unobserved intermediate, once we condition on the observed one, has a smaller impact on the outcome than the observed one had. This can be seen as a weakened version of principal ignorability as the principal ignorability assumption assumes that there is no effect of the unobserved intermediate, once we condition on the observed one. This is formalized in the following assumption.
@@
$\begin{split}&\Bigg{[}-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}},-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg{]}\cup\\ &\qquad\Bigg{[}\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg{]},\end{split}$
@@
Note that the result depends on $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$, ${(\text{Var}\{Y(0)\mid S(0)\})^{2}}\break/{\text{Var}(Y(0))}\geq{(\text{Var}\{Y(1)\mid S(1)\})^{2}}/{\text{Var}(Y(1))}$, and $\beta_{00}>0$ all being true. We should emphasize that similar bounds can be derived under any combination of directions for these inequalities, but we focus on one such choice for simplicity. By comparing the results in Propositions 1 and 3, it is clear that Assumption 7 reduces the length of the partial identification region substantially. For $\beta_{01}$ the size of the region is reduced by a factor of $|\text{Cor}(Y(0),S(0))|$, which reflects the strength of the association between the outcome and the observed intermediate within the control $(T=0)$ group. When $|\text{Cor}(Y(0),S(0))|$ is very small, the benefits brought by Assumption 7 are large since it prevents assigning overly large values to the effect of the missing intermediate.
@@ 6 Simulations and applications
Here, we examine the theoretical results obtained in the previous sections as well as the practical performance of our weakened identification assumptions for estimating principal causal effects. In the first set of simulation studies, we treat $\rho$ as known, and focus on partial identification of principal causal effects of interest under varying assumptions. Then, we explore the identification of $\rho$ under varying assumptions and sample sizes. Throughout, we do not include covariates $\boldsymbol{X}$, but similar results would be obtained if covariates were additionally adjusted for. Under both scenarios, Gibbs sampling is used, with a single MCMC chain run for a total of 25,000 iterations, where the first 5,000 are burn-in, and a thinning interval of 30 is applied. The details of the Gibbs sampler are included in Appendix A.
6.1 Performance under a known $\rho$
@@
$\begin{split}Y_{i}(0)\mid\boldsymbol{U}_{i}&\sim N(-0.5+11.5S(0),14^{2}),\\ Y_{i}(1)\mid\boldsymbol{U}_{i}&\sim N(-0.5+11.5S(0)+96S(1),14^{2}),\\ \boldsymbol{U}_{i}&\sim N((0.89,0.70)^{T},\boldsymbol{\Sigma}_{S}),\end{split}$ (4)
@@
True value Mean ECR Width Mean ECR Width Mean ECR Width
$\boldsymbol{u}=(0.89,0.18)$ 17 17 0.99 43 16 1.00 25 23 0.97 32
$\boldsymbol{u}=(0.89,0.35)$ 34 34 0.99 13 34 1.00 9 34 0.98 12
$\boldsymbol{u}=(0.89,0.52)$ 50 50 0.99 43 51 1.00 25 45 0.98 32
@@
6.2 Identifying $\rho$ under different constraints
Now we use a related data generating process to further explore the ability to identify $\rho$ under varying assumptions. As discussed previously, one plausible identification assumption for $\rho$ is $\beta_{01}=0$ and $\beta_{00}=\beta_{10}$. This guarantees the identification of $\rho$ when we also assume $\rho>0$. To investigate both identification and posterior consistency, we vary $n$ within {300, 600, 1200} and we generate one dataset for each sample size. For each dataset, we run MCMC under no constraints, one constraint ($\beta_{01}=0$), and two constraints ($\beta_{01}=0$, $\beta_{00}=\beta_{10}$). To avoid the sign issue, we also constrain $\rho$ to be in $(0.00,0.95)$, where setting $\rho<0.95$ prevents MCMC sampling from getting trapped around the extreme value $\rho=1$.
@@ 6.2 Identifying $\rho$ under different constraints
$\begin{split}Y_{i}(0)\mid\boldsymbol{U}_{i}&\sim N(0.9+1.2S(0),0.5^{2}),\\ Y_{i}(1)\mid\boldsymbol{U}_{i}&\sim N(0.5+1.2S(0)+1.2S(1),0.5^{2}),\\ \boldsymbol{U}_{i}&\sim N((0.3,0.5)^{T},\boldsymbol{\Sigma}_{S}),\end{split}$ (5)
where $\boldsymbol{\Sigma}_{S}=\begin{pmatrix}1&0.75\\ 0.75&1\end{pmatrix}$. The same noninformative conjugate priors are used and we additionally place a flat prior on the correlation between potential intermediates, $\rho\sim U(0,0.95)$.
@@
The posterior distributions of $\rho$ across all three scenarios and sample sizes are shown in Figure 1. Regardless of sample size, there is a notable spike around the true value of $\rho=0.75$ when we place two constraints on the model parameters. There appear to be mild spikes close to 1 when there are no constraints or one constraint placed on the model, but this is simply due to poor mixing as the MCMC sampler gets stuck at $\rho$ values close to 1. When we alternatively assume $\rho\sim U(0,0.9)$, these spikes are reduced. Importantly, as $n$ increases, their height does not increase, whereas the height of spike under two constraints increases significantly, which points towards consistency of the correlation parameter. In Appendix C, we include more plots of the estimated posterior variance for more sample sizes, which show that asymptotically, the posterior variance is $O(n^{-1})$.
@@
Figure 1: The posterior distributions of $\rho$ corresponding to $n=$300, 600, 1200 under three different amounts of constraints on the model parameters.

### v3

@@ 1 Introduction
Another line of research aims to incorporate parametric assumptions to obtain identification when nonparametric identification is difficult or impossible to obtain. This is particularly the case with continuous intermediate variables, which we address in this manuscript, as standard assumptions such as monotonicity and exclusion restrictions are not sufficient for nonparametric identification. Recent works studying nonparametric identification in this more difficult scenario have made the principal ignorability assumption, along with the strong parametric assumption that the joint distribution of the two potential intermediates is governed by a known copula with a known correlation parameter [21, 28, 42]. Results are then estimated across a range of plausible correlation parameters to see if results are robust to this choice. Other works have attempted to estimate this unknown correlation parameter by jointly modeling the potential intermediates and observed outcomes simultaneously [37, 4]. The idea behind this strategy is that one cannot directly estimate the association between the two potential intermediates because they are never jointly observed, but this information can be obtained from a correctly specified outcome regression model. Related research aims to jointly model the distribution of the potential intermediates and observed outcome using either parametric [22] or semiparametric models [24, 25], and assume the correlation between the two potential intermediates is known. Many of these model-based approaches do not make the principal ignorability assumption, and they allow the observed outcome to depend on the values of both potential intermediate values.
@@ 2.2 Review of identification assumptions in PS analysis
$0<P(T=1\mid\boldsymbol{X}=\boldsymbol{x})<1$ for all $\boldsymbol{x}$.
@@ 2.3 Binary intermediates
While plausible in certain settings, such as those with randomized treatment assignment and noncompliance, these are frequently not likely to hold, as there are many settings where a direct effect of treatment is plausible. In such settings, one can instead use the following assumption.
@@
2.4 Continuous intermediates
[…]
@@ 3.1 A simple parametric model
[…]
@@
[…]
@@ 3.2 Identification of the association parameter of the principal strata
where $\bar{T}=\sumop\displaylimits\limits_{i=1}^{n}T_{i}/n$, $\zeta^{*}_{0}-{\psi^{*}_{0}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}$, and $\zeta_{1}^{*}-{\psi^{*}_{1}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}$.
@@ 4.1 Partial identification and principal causal effects
[…]
@@ 5 Weaker alternative assumptions of principal ignorability
For $\beta_{00}$ and $\beta_{11}$ with signs different from those assumed in Proposition 2, the corresponding partial identification regions for $\beta_{01}$ and $\beta_{10}$ can be obtained by flipping the signs to match those of $\beta_{00}$ and $\beta_{11}$, respectively. This result also requires us to know the sign of $\beta_{tt}$ for $t=0,1$. If this is not known a priori, then one can regress the observed $Y$ on the observed $S$ and $\boldsymbol{X}$, given $T=t$, to determine the sign for $\beta_{tt}$. This approach is not guaranteed to correctly identify the sign, but will only fail to do so in certain extreme situations, such as when $\rho$ is close to 1-1. This cuts the partial identification region in half, providing large benefits in the uncertainty in the resulting causal estimates. While these benefits may be large, it is not possible in many situations to know this assumption is true a priori. Now we present a second, weaker assumption, which should hold true in most applications. This second assumption effectively assumes that the unobserved intermediate, once we condition on the observed one, has a smaller impact on the outcome than the observed one had. This can be seen as a weakened version of principal ignorability as the principal ignorability assumption assumes that there is no effect of the unobserved intermediate, once we condition on the observed one. This is formalized in the following assumption.
@@
$\begin{split}&\Bigg[-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}},-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg]\cup\\ &\hskip 17.00024pt\Bigg[\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg],\end{split}$
@@
Note that the result depends on $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$, ${(\text{Var}\{Y(0)\mid S(0)\})^{2}}\penalty-10000/{\text{Var}(Y(0))}\geq{(\text{Var}\{Y(1)\mid S(1)\})^{2}}/{\text{Var}(Y(1))}$, and $\beta_{00}>0$ all being true. We should emphasize that similar bounds can be derived under any combination of directions for these inequalities, but we focus on one such choice for simplicity. By comparing the results in Propositions 1 and 3, it is clear that Assumption 7 reduces the length of the partial identification region substantially. For $\beta_{01}$ the size of the region is reduced by a factor of $|\text{Cor}(Y(0),S(0))|$, which reflects the strength of the association between the outcome and the observed intermediate within the control $(T=0)$ group. When $|\text{Cor}(Y(0),S(0))|$ is very small, the benefits brought by Assumption 7 are large since it prevents assigning overly large values to the effect of the missing intermediate.
@@ 6 Simulations and applications
6.1 Simulations
Here, we examine the theoretical results obtained in the previous sections as well as the practical performance of our weakened identification assumptions for estimating principal causal effects. In the first set of simulation studies, we treat $\rho$ as known, and focus on partial identification of principal causal effects of interest under varying assumptions. Then, we explore the identification of $\rho$ under varying assumptions and sample sizes. Throughout, we do not include covariates $\boldsymbol{X}$, but similar results would be obtained if covariates were additionally adjusted for. Under both scenarios, Gibbs sampling is used, with a single MCMC chain run for a total of 25,000 iterations, where the first 5,000 are burn-in, and a thinning interval of 30 is applied. The details of the Gibbs sampler are included in Appendix B.
6.1.1 Performance under a known $\rho$
@@
[…]
@@
$\boldsymbol{u}$ Truth Mean ECR Width Mean ECR Width Mean ECR Width
$(0.89,0.18)$ 17 17 0.99 43 16 1.00 25 23 0.97 32
$(0.89,0.35)$ 34 34 0.99 13 34 1.00 9 34 0.98 12
$(0.89,0.52)$ 50 50 0.99 43 51 1.00 25 45 0.98 32
@@
6.1.2 Identifying $\rho$ under different constraints
Now we use a related data generating process to further explore the ability to identify $\rho$ under varying assumptions. As discussed previously, one plausible identification assumption for $\rho$ is $\beta_{01}=0$ and $\beta_{00}=\beta_{10}$. This guarantees the identification of $\rho$ when we also assume $\rho>0$. To investigate both identification and posterior consistency, we vary $n\in\{300,600,1200\}$ and we generate one dataset for each sample size. For each dataset, we run MCMC under no constraints, one constraint ($\beta_{01}=0$), and two constraints ($\beta_{01}=0$, $\beta_{00}=\beta_{10}$). To avoid the sign issue, we also constrain $\rho$ to be in $(0.00,0.95)$, where setting $\rho<0.95$ prevents MCMC sampling from getting trapped around the extreme value $\rho=1$.
@@ 6.1.2 Identifying $\rho$ under different constraints
[…]
@@
The posterior distributions of $\rho$ across all three scenarios and sample sizes are shown in Figure 1. Regardless of sample size, there is a notable spike around the true value of $\rho=0.75$ when we place two constraints on the model parameters. There appear to be mild spikes close to 1 when there are no constraints or one constraint placed on the model, but this is simply due to poor mixing as the MCMC sampler gets stuck at $\rho$ values close to 1. When we alternatively assume $\rho\sim U(0,0.9)$, these spikes are reduced. Importantly, as $n$ increases, their height does not increase, whereas the height of spike under two constraints increases significantly, which points towards consistency of the correlation parameter. In Appendix D, we include more plots of the estimated posterior variance for more sample sizes, which show that asymptotically, the posterior variance is $O(n^{-1})$.
@@
Figure 1: The posterior distributions of $\rho$ corresponding to n=n=300, 600, 1200 under three different amounts of constraints on the model parameters.
6.2 Analysis of ACTG trial data
The ACTG 175 data set ([15]) is available in the R package speff2trial and was collected from a randomized clinical trial with the purpose of comparing monotherapy with zidovudine or didanosine with combination therapy with zidovudine and didanosine or zidovudine and zalcitabine in adults infected with the human immunodeficiency virus type I whose CD4 T cell counts were between 200 and 500 per cubic millimeter. This dataset was investigated by [6] and [42], with a focus on whether a short-term endpoint can serve as a valid surrogate for a long-term endpoint. Here, we explore the same question for this dataset, though we do not assume principal ignorability holds as in the analysis of [42], and we explore the extent to which our various proposed assumptions can provide informative partial identification regions. We consider the short-term endpoint, CD4 count at $20\pm 5$ weeks, as the intermediate variable $S$, and the long-term endpoint, CD4 count at $96\pm 5$ weeks, as the outcome $Y$. The treatment $T$ is 0 if the patient is treated with zidovudine only and is 1 if the patient is treated with zidovudine + didanosine. Covariates $\boldsymbol{X}$ include age, weight, hemophilia, homosexual activity, history of intravenous drug use, Karnofsky score, non-zidovudine antiretroviral therapy prior to initiation of study treatment, zidovudine use in the 30 days prior to treatment initiation, race, gender, antiretroviral history, and a symptomatic indicator. We apply Model (1) throughout with a fixed value of $\rho=0.5$. Note that this model does not assume principal ignorability, which we show in Section 4 can lead to substantially wider credible intervals for the principal causal effects of interest. For this reason, we also explore incorporating both the dominant effect assumption and the same sign assumptions proposed in Section 5, which leads to three different approaches in total as we also consider the case where no assumptions are incorporated. Throughout we consider a range of principal strata that can be broken into two types. The first looks at subjects for whom $S(0)=S(1)=s$ and we vary the value of $s$. The second set of principal strata fixes the value of $S(0)=340$ and then varies $S(1)$ to see how the causal effect for the final endpoint depends on the effect in the surrogate endpoint.
The results can be found in Figure 2, which shows both the posterior mean and corresponding 95% credible interval for all estimands of interest across the three proposed approaches. One of the most apparent findings from the results is that inference without principal ignorability and no additional assumptions (first column of Figure 2) is less informative with wide credible intervals, but the two proposed assumptions (2nd and 3rd columns) can substantially shrink the size of the credible intervals, particularly for the same sign assumption. The results under both the dominant effect and same sign assumptions also paint a fairly clear picture about the performance of $S$ as a surrogate endpoint. When $S(0)=S(1)=s$, there are slightly positive estimates with credible intervals that generally contain zero, though the intervals under the same sign assumption are strictly positive for certain values of $s$. This suggests that if there is no effect of the treatment by the surrogate endpoint at 20 weeks, then there likely won’t be a big effect of the treatment by the terminal endpoint at 96 weeks, though the results are suggestive that there is still some beneficial effect of treatment even when $S(0)=S(1)$. This indicates that it may take time for the full effect of treatment to realize, though this effect is not substantial. The results from when $S(0)\neq S(1)$ also show that the surrogate endpoint is a valid surrogate in the sense that when S(1)S(0)S(1)-S(0) is larger, the treatment effect for the final endpoint described by the magnitude of Y(1)Y(0)Y(1)-Y(0) is also large. These results are similar to those obtained in recent analyses assuming principal ignorability [42]. This shows the benefit of our proposed assumptions, which are arguably weaker than principal ignorability, but are still able to provide meaningful insights on principal causal effects. If principal ignorability holds, we can still obtain similar results without greatly increasing the widths of the credible intervals, while if principal ignorability fails, our results should be more robust.
Figure 2: The first row shows PCEs and corresponding 95% credible intervals for the principal strata defined by $S(0)=S(1)=s$ with varying $s$. The second row corresponds to the principal strata with $S(0)=340$ and increasing $S(1)$. Note that $S(0)$ and $S(1)$ are scaled by $10^{2}$ in the plot. From left to right, the columns correspond to different assumptions: no constraints, dominant effect, and same sign, respectively. The dashed line in the CI plots implies that the CI is truncated, and the arrow indicates which side is truncated.